OpenStudy (anonymous):

Simplify, log1/2 8. (1/2 is the base) Thanks.

OpenStudy (whpalmer4):

\[\log_{1/2} 8 = n\] where \[(\frac{1}{2})^n = 8\]

OpenStudy (whpalmer4):

Hint: \[(\frac{a}{b})^n = (\frac{b}{a})^{-n}\]

OpenStudy (anonymous):

-3?

OpenStudy (whpalmer4):

Yes! :-)

OpenStudy (anonymous):

Thank you !!

OpenStudy (whpalmer4):

You're welcome.